OpenStudy (anonymous):

Two equations are given below: m + 5n = 20 m = n − 4 What is the solution to the set of equations in the form (m, n)?

pooja195 (pooja195):

use substituion We know the m=n-4 right? So we plug that in place of m in equation 1 n-4+5n=20 solve for n please :)

OpenStudy (anonymous):

n=4? @pooja195

pooja195 (pooja195):

good now that we know n=4 we plug that into equation 2 into the place of n 4-4 =?

OpenStudy (anonymous):

m=0? @pooja195

pooja195 (pooja195):

yes :) so our answer is?

OpenStudy (anonymous):

(0,4) :D @pooja195

pooja195 (pooja195):

yes! :D

OpenStudy (anonymous):

Thank you so much, wow. That was easy. @pooja195

pooja195 (pooja195):

Glad i could help and im happy u think its easy!: )